OpenStudy (voltage):

WILL FAN AND MEDAL HEEEEEELPPP!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

What is the question?

OpenStudy (voltage):

Rewrite 4 − 5 using the additive inverse and display the new expression on a number line. 4 + 5 A horizontal number line is shown with labels from 0 to 16. A blue arrow begins at 0 and ends at 4. Another blue arrow begins at 4 and ends at 9. 4 − (−5) A horizontal number line is shown with labels from 0 to 16. A blue arrow begins at 0 and ends at 4. Another blue arrow begins at 4 and ends at 9. 5 − 4 A horizontal number line is shown with labels from negative 8 to positive 8. A blue arrow begins at 0 and ends at 5. A red arrow begins at 5 and ends at 1. 4 + (−5) A horizontal number line is shown with labels from negative 8 to positive 8. A blue arrow begins at 0 and goes to 4. A red arrow begins at 4 and goes to negative 1. Question 2 (Multiple Choice Worth 5 points) Use absolute value to express the distance between −10 and 16 on the number line. |−10 − 16| = −6 |−10 − 16| = 6 |−10 − 16| = −26 |−10 − 16| = 26 Question 3 (Multiple Choice Worth 5 points) What number should be added to 17 to make the sum 0? 0 –16 17 –17 Question 4 (Multiple Choice Worth 5 points) Simplify: Negative three and one-half minus negative twelve and one-fourth. Negative fifteen and three-fourths Fifteen and three-fourths Negative eight and three-fourths Eight and three-fourths Question 5 (Essay Worth 5 points) Determine the answer to 3 + (−5) and explain the steps using a number line. Question 6 (Essay Worth 5 points) Andrew is showing his work in simplifying −4.5 + 4.2 + 5.6 − 7.3. Identify any errors in his work or in his reasoning. Write feedback to Andrew explaining the mistake he made and how to correct his work. Step Work Reasoning −4.5 + 4.2 + 5.6 − 7.3 Original problem 1 −4.5 + 4.2 + 5.6 + (−7.3) Additive inverse 2 −4.5 + (−7.3) + 4.2 + 5.6 Commutative property 3 (−4.5 + (−7.3)) + (4.2 + 5.6) Associative property 4 −11.8 + 9.8 Simplify 5 −21.6 Simplify @shadyvirus

OpenStudy (anonymous):

For question number 2, what you need to do is subtract 16 from -10 and find the absolute of value of the answer.

OpenStudy (voltage):

alright

OpenStudy (anonymous):

For question 3 you are finding the additive inverse of 17, which is what number can you add to 17 to make it zero. So basically you would take a positive number and add the negative of that number, which is the additive inverse. Ex. 5+(-5)=0 with -5 being the additive inverse.

OpenStudy (anonymous):

For question 4, first you need to write out the equation. \[-3\frac{ 1 }{ 2 } - \left( -12\frac{ 1 }{ 4 } \right)\] From there you need to turn the fractions into a mixed number. To do that, you multiply the denominator by the whole number, then add the numerator. \[-3\frac{ 1 }{ 2 } = -7/2\] and \[-12\frac{ 1 }{ 4 } = -\frac{ 48 }{ 4 }\] Your problem will read -7/2 - (-48/4), change this to -7/2 + 48/4. Now, we need to find a common denominator, to do this, we will change -7/2 to -14/4. -14 + 49 = 35. 35/4 = 8.75 or 8 3/4 which also reads as eight and three-fourths.

OpenStudy (anonymous):

Unfortunately, I cannot help you with the rest.